# Quantum Blackwell’s Ordering and Differential Privacy

### v1

@@ VIII-A Formulation of $(\varepsilon,\delta)$-LDP classical channels
A classical channel $\mathcal{K}$ is $(\varepsilon,\delta)$-LDP, if and only if any pair of distributions $P_{X},Q_{X}\in\mathcal{P}(\mathcal{X})$ satisfies the relation
@@
VIII-B Contraction Coefficient for Divergences under Private Classical/Quantum Learning Algorithms
For any classical divergence $\mathbb{D}$, from Definition 9, we know that for any pair of probability distributions $P,Q\in\mathcal{P}(\mathcal{X})$ (where $\mathcal{X}$ is some finite set) and a classical channel $\mathcal{K}:\mathcal{X}\to\mathcal{Y}$ (where $\mathcal{Y}$ is another finite set), $\mathbb{D}(\mathcal{K}(P_{1})\|\mathcal{K}(P_{2}))\leq\mathbb{D}(P_{1}\|P_{2})$, where for each $i=1,2$, $\mathcal{K}(P_{i})$ is the marginal output distribution of $\mathcal{K}$ corresponding to $P_{i}$. However, this inequality is not always strict. Thus, we are interested in how much the channel $\mathcal{K}$ shrinks the KL divergence for the privacy of the respondent. To clarify this, we study the largest constant $\eta^{(c),\mathcal{K}}_{\mathbb{D}}$ satisfying the condition $\mathbb{D}(\mathcal{K}(P_{1})\|\mathcal{K}(P_{2}))\leq\eta^{(c),\mathcal{K}}_{\mathbb{D}}\mathbb{D}(P_{1}\|P_{2})$ for all pairs of distributions $P_{1},P_{2}\in\mathcal{P}(\mathcal{X})$. The quantity $\eta^{(c),\mathcal{K}}_{\mathbb{D}}$ is called the contraction coefficient of the channel $\mathcal{K}$ with respect to the divergence $\mathbb{D}$.
@@ VIII-B Contraction Coefficient for Divergences under Private Classical/Quantum Learning Algorithms
Hirche et al. in [32] have studied the following bound on $\eta^{\mathcal{N}}_{E_{\gamma}}$ for any CP-TP some fixed $\gamma\geq 1$.
Proposition 3 ([32, Lemma II.4]).
@@
In the following lemma, we now generalize the above result for the contraction coefficient of quantum hockey stick divergence $E_{\gamma}(\cdot\|\cdot)$ for any $\gamma\geq 1$ under $(\varepsilon,\delta)$-LDP quantum mechanisms for any $\varepsilon\geq 0$ and $\delta\in[0,1]$.
@@
$\eta_{E_{\gamma}}^{\varepsilon,\delta}$ $\overset{a}{\geq}1-\gamma(1-\eta^{\varepsilon,\delta}_{E_{1}})$
$\overset{b}{=}1-2\gamma\frac{(1-\delta)}{e^{\varepsilon}+1}$
@@
$\overset{c}{\geq}\frac{e^{\varepsilon}+\delta-\gamma(1-\delta)}{e^{\varepsilon}+1}+\frac{(1-\gamma)}{e^{\varepsilon}+1}$
@@
where $a$ follows from Proposition 3, $b$ follows from Proposition 4, $c$ follows because $\gamma\geq 1.$
@@
    $\eta_{E_{\gamma}}^{\varepsilon,\delta}$ $\overset{a}{=}\sup_{\begin{subarray}{c}\mathcal{N}\in\mathfrak{P}_{\varepsilon,\delta},\\ |\psi\rangle\perp|\phi\rangle\end{subarray}}E_{\gamma}(\mathcal{N}(|\psi\rangle\langle\psi|)\|\mathcal{N}(|\phi\rangle\langle\phi|))$
    $\overset{b}{\leq}E_{\gamma}(\rho_{(\varepsilon,\delta)}\|\sigma_{(\varepsilon,\delta)})$
    $\overset{c}{=}\frac{e^{\varepsilon}-\gamma+\delta(\gamma+1)}{e^{\varepsilon}+1},$ (100)
@@
    where in $a$ $|\psi\rangle,|\phi\rangle\in\mathcal{H}_{A}$ are two orthogonal quantum states and the inequality follows from [32, Theorem II.2], $b$ follows because $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})$ is the weakest (most informative) $(\varepsilon,\delta)$-DP pair of quantum states as mentioned in eqs.˜51 and 53 and $c$ follows from Claim 1 below.
@@
    where $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})$ is the weakest (most informative) $(\varepsilon,\delta)$-DP pair of quantum states as mentioned in eqs.˜51 and 53.
@@
In the quantum setting, Nuradha et al in [20, Proposition 3] obtain an upper-bound on quantum relative entropy under $\varepsilon$-LDP CP-TP maps via an integral representation of quantum relative entropy (Fact 6) in terms of quantum hockey stick divergence. Moreover, the upper-bound obtained in [20, Proposition 3] is tighter than the upper-bound obtained in $(i)$ of Corollary 4.
@@
To resolve this issue, in the subsection below, we come up with a technique which we call truncation. This technique allows us to distill pure DP from non-pure DP.
VIII-C upper-bound on the relative entropy of $(\varepsilon,\delta)$-LDP classical channels via the integral representation
@@
See Appendix -I for the proof.
Alternatively, we can obtain a different upper-bound on $D(\mathcal{K}(P_{X})\|\mathcal{K}(Q_{X}))$ by directly applying the continuity bound for relative entropy from (ii) of Lemma 8 and substituting $\varepsilon\leftarrow\left(\varepsilon+\log\frac{1}{1-\delta}\right)$ in Corollary 4.
@@
It is important to note that the bound obtained in Theorem 6 is tighter than the bound in Theorem 7. This is because the bound in Theorem 7 is obtained by first approximating the $(\varepsilon,\delta)$-DP pair with an $(\varepsilon^{\prime},0)$-DP pair, where $\varepsilon^{\prime}=\varepsilon+\log\frac{1}{1-\delta}$, and then applying the known bounds for pure DP. This approximation introduces looseness, particularly in the leading term, which becomes $\mathcal{O}((\varepsilon+\delta)\tanh(\varepsilon+\delta))$. In contrast, Theorem 6 uses a more direct approach via the integral representation of the KL divergence, resulting in a leading term of $\mathcal{O}(\varepsilon\tanh(\varepsilon)){\left\lVert P_{X}-Q_{X}\right\rVert}_{1}$. For small $\varepsilon$, the bound in Theorem 6 is therefore significantly tighter.
IX Conclusion
In this work, inspired by [7], we develop a framework for studying quantum differential privacy from the perspective of hypothesis testing and Blackwell’s ordering [9]. We provided a characterization of quantum $(\varepsilon,\delta)$-differential privacy in terms of quantum hypothesis testing divergences, and identified the most informative (in the Blackwell sense) pair of quantum states. We use this framework for studying the stability of differentially private quantum learning algorithms. Our stability result also generalizes (in the sense that $\delta>0$) the existing bound in the classical settings.
We also study the problem of quantum privatized parameter estimation, where the trade-off between privacy and statistical utility is characterized via the quantum Fisher information. We derived explicit expressions for the maximal Fisher information achievable under quantum $(\varepsilon,\delta)$-DP constraints, thereby quantifying the fundamental limits of parameter estimation in the presence of quantum privacy mechanisms.
@@ Acknowledgments
The work of NAW was supported in part by MTR/2022/000814. The work of MH was supported in part by the National Natural Science Foundation of China under Grant 62171212 and in part by the General Research and Development Projects of 1+1+1 CUHK-CUHK(SZ)-GDST Joint Collaboration Fund under Grant GRDP2025-022.

### v3

@@ VI-A Formulation of $(\varepsilon,\delta)$-LDP classical channels
A classical channel $\mathcal{K}$ is $(\varepsilon,\delta)$-LDP, if and only if any pair of distributions $P,Q\in\mathcal{P}(\mathcal{X})$ satisfies the relation
@@
VI-B Contraction Coefficient for Divergences under Private Classical/Quantum Learning Algorithms
For any classical divergence $\mathbb{D}$, from Definition 6, we know that for any pair of probability distributions $P_{1},P_{2}\in\mathcal{P}(\mathcal{X})$ (where $\mathcal{X}$ is some finite set) and a classical channel $\mathcal{K}:\mathcal{X}\to\mathcal{Y}$ (where $\mathcal{Y}$ is another finite set), $\mathbb{D}(\mathcal{K}(P_{1})\|\mathcal{K}(P_{2}))\leq\mathbb{D}(P_{1}\|P_{2})$, where for each $i=1,2$, $\mathcal{K}(P_{i})$ is the marginal output distribution of $\mathcal{K}$ corresponding to $P_{i}$. However, this inequality is not always strict. Thus, we are interested in how much the channel $\mathcal{K}$ shrinks the KL divergence for the privacy of the respondent. To clarify this, we study the largest constant $\eta^{(c),\mathcal{K}}_{\mathbb{D}}$ satisfying the condition $\mathbb{D}(\mathcal{K}(P_{1})\|\mathcal{K}(P_{2}))\leq\eta^{(c),\mathcal{K}}_{\mathbb{D}}\mathbb{D}(P_{1}\|P_{2})$ for all pairs of distributions $P_{1},P_{2}\in\mathcal{P}(\mathcal{X})$. The quantity $\eta^{(c),\mathcal{K}}_{\mathbb{D}}$ is called the contraction coefficient of the channel $\mathcal{K}$ with respect to the divergence $\mathbb{D}$.
@@ VI-B Contraction Coefficient for Divergences under Private Classical/Quantum Learning Algorithms
Hirche et al. in [18] studied the following bound on $\eta^{\mathcal{N}}_{E_{\gamma}}$ for any CP-TP some fixed $\gamma\geq 1$.
Proposition 2 ([18, Lemma II.4]).
@@
In the following lemma, we now generalize the above result for the contraction coefficient of quantum hockey stick divergence $E_{\gamma}(\cdot\|\cdot)$ for any $\gamma\geq 1$ under $(\varepsilon,\delta)$-LDP CP-TP maps for any $\varepsilon\geq 0$ and $\delta\in[0,1]$.
@@
$\eta_{E_{\gamma}}^{\varepsilon,\delta}$ $\overset{(a)}{\geq}1-\gamma(1-\eta^{\varepsilon,\delta}_{E_{1}})$
$\overset{(b)}{=}1-2\gamma\frac{(1-\delta)}{e^{\varepsilon}+1}$
@@
$\overset{(c)}{\geq}\frac{e^{\varepsilon}+\delta-\gamma(1-\delta)}{e^{\varepsilon}+1}+\frac{(1-\gamma)}{e^{\varepsilon}+1}$
@@
where $(a)$ follows from Proposition 2, $(b)$ follows from Proposition 3, $(c)$ follows because $\gamma\geq 1.$
@@
    $\eta_{E_{\gamma}}^{\varepsilon,\delta}$ $\overset{(a)}{=}\sup_{\begin{subarray}{c}\mathcal{N}\in\mathfrak{P}_{\varepsilon,\delta},\\ |\psi\rangle\perp|\phi\rangle\end{subarray}}E_{\gamma}(\mathcal{N}(|\psi\rangle\langle\psi|)\|\mathcal{N}(|\phi\rangle\langle\phi|))$
    $\overset{(b)}{\leq}E_{\gamma}(\rho_{(\varepsilon,\delta)}\|\sigma_{(\varepsilon,\delta)})$
    $\overset{(c)}{=}\frac{e^{\varepsilon}-\gamma+\delta(\gamma+1)}{e^{\varepsilon}+1},$ (104)
@@
    where in $(a)$, $|\psi\rangle,|\phi\rangle\in\mathcal{H}_{A}$ are two orthogonal quantum states and the inequality follows from [18, Theorem II.2], $(b)$ follows from Theorem 2 and $(c)$ follows from Claim 1 below.
@@
    where $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})$ is the weakest (most informative) $(\varepsilon,\delta)$-DP pair of quantum states as mentioned in eqs.˜47 and 49.
@@
In the quantum setting, Nuradha et al in [26, Proposition 3] obtained an upper-bound on the quantum relative entropy under $\varepsilon$-LDP CP-TP maps via an integral representation of quantum relative entropy (Fact 6) in terms of quantum hockey stick divergence. Moreover, the upper-bound obtained in [26, Proposition 3] is tighter than the upper-bound obtained in $(i)$ of Corollary 3.
@@
In the subsection below, we resolve this issue in classical setting, by coming up with a technique which we call truncation. This technique allows us to distill pure DP from non-pure DP.
VI-C Upper-bound on the relative entropy of $(\varepsilon,\delta)$-LDP classical channels via the integral representation
@@
See Appendix -F for the proof.
Alternatively, in the theorem below, we obtain a different upper-bound on $D(\mathcal{K}(P_{X})\|\mathcal{K}(Q_{X}))$ by directly applying the continuity bound for relative entropy from $(ii)$ of Lemma 6 and substituting $\varepsilon\leftarrow\left(\varepsilon+\log\frac{1}{1-\delta}\right)$ in Corollary 3.
@@
It is important to note that the bound obtained in Theorem 5 is tighter than the bound in Theorem 6. This is because the bound in Theorem 6 is obtained by first approximating the $(\varepsilon,\delta)$-DP pair with an $\varepsilon^{\prime}$-DP pair, where $\varepsilon^{\prime}=\varepsilon+\log\frac{1}{1-\delta}$, and then applying the known bounds for pure DP. This approximation introduces looseness, particularly in the leading term, which becomes $\mathcal{O}((\varepsilon+\delta)\tanh(\varepsilon+\delta))$. In contrast, Theorem 5 uses a more direct approach via the integral representation of the KL divergence, resulting in a leading term of $\mathcal{O}(\varepsilon\tanh(\varepsilon)){\left\lVert P_{X}-Q_{X}\right\rVert}_{1}$. For small $\varepsilon$, the bound in Theorem 5 is therefore significantly tighter.
VII Conclusion
In this work, motivated by Blackwell’s order of informativeness [4, 3] between two statistical experiments, we define an order of informativeness in the quantum setting, which is based on the hypothesis testing divergence dominance. In the classical case, Blackwell’s order implies that there always exists a Markov kernel that maps the more informative statistical experiment to the lesser informative statistical experiment. However, in the quantum case, a CP-TP map may not always exist [23] from a more informative pair of quantum states to the lesser informative pair of quantum states. Despite the lack of existence of such CP-TP maps, we still upper-bound the $f$-divergence of lesser informative pair of quantum states by that of a more informative pair of quantum states. This approach allows us to fully characterize $(\varepsilon,\delta)$-QDP mechanisms.
We use this ordering to study the problem of quantum privatized hypothesis testing and quantum privatized parameter estimation, where the trade-off between privacy and statistical utility is characterized via the quantum hypothesis testing divergence and the quantum Fisher information, respectively. We derived explicit expressions for the maximal Fisher information achievable under quantum $(\varepsilon,\delta)$-DP constraints, thereby quantifying the fundamental limits of parameter estimation in the presence of quantum privacy mechanisms.
@@ Acknowledgments
We are grateful to Henrik Wilming for pointing out an error in the earlier version of this manuscript. His comments have greatly helped us to make the technical contributions of this manuscript more rigorous. We thank Francesco Buscemi for pointing out several useful references and also thank Mark Wilde for his comments on the earlier version.
The work of NAW was supported by MTR/2022/000814. The work of AD was supported by TCS Research Scholar Program (Cycle 19). The work of MH was supported in part by the General R&D Projects of 1+1+1 CUHK-CUHK(SZ)-GDST Joint Collaboration Fund (Grant No. GRDP2025-022), the Guangdong Provincial Quantum Science Strategic Initiative (Grant No. GDZX2505003), the Shenzhen International Quantum Academy (Grant No. SIQA2025KFKT07), and the National Natural Science Foundation of China under Grant 62171212.